# Fitness Landscape of Large Language Model-Assisted Automated Algorithm Search

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated significant potential in algorithm design. However, when integrated into search frameworks for iterative algorithm search, the underlying fitness landscape—critical for understanding search behaviour—remains underexplored. In this paper, we illustrate and analyze the fitness landscape of LLM-assisted Algorithm Search (LAS) using a graph-based approach, where nodes represent algorithms and edges denote transitions between them. We conduct extensive evaluations across six algorithm design tasks and six commonly used LLMs. Our findings reveal that LAS landscapes are highly multimodal and rugged, particularly in combinatorial optimization tasks, with distinct structural variations across tasks and LLMs. For instance, heuristic design tasks exhibit dense clusters of high-performing algorithms, while symbolic regression tasks show sparse, scattered distributions. Additionally, we demonstrate how population size influences exploration-exploitation trade-offs and the evolving trajectory of elite algorithms. These insights not only advance our understanding of LAS landscapes but also provide practical guidance for designing more effective LAS methods.
@@ I Introduction
Large Language Models (LLMs) assisted iterative search methods have been used for automated algorithm design with great success [1] [2] [3] [4] [5] [6]. These methods cast an algorithm design task as an optimization problem for finding an optimal algorithm in an algorithm space. Leveraging the LLM capabilities in language comprehension, reasoning and code generation, they are able to conduct effective and efficient search in an algorithm space [7]. However, there is lack of good understanding of why and when these methods work or not. Traditional math analysis can be hardly used to study these methods due to the complexity of the algorithm space [8] and the black-box nature of LLM [9]. Fitness landscape analysis [10, 11, 12, 13, 14, 15] is a widely used methodology for investigating behaviors of search methods, particularly, in the area of heuristics. Very recently, [16] has made a very first attempt to study the landscape of LLM-assisted algorithm design methods. They introduced abstract syntax tree (AST)-based code evolution graphs (CEGs) and investigated the relationship between candidate algorithms (i.e., an element in the algorithm space) and their characters in feature space. However, their work does not discuss the fitness of a candidate algorithm  [^1] For the sake of clarity, we call an algorithm in the algorithm space a candidate algorithm.
@@
In this paper, we focus on the fitness landscape of LLM-assisted Algorithm Search (LAS). Leveraging a graph-based approach, we try to visualize and analyze the fitness landscape, providing deeper insights into the search dynamics of LAS. Specifically, we try to answer the following questions:
@@
  • Q1: What does the fitness landscape of LLM-driven algorithm search look like, and what features (e.g., plateaus, ruggedness) characterize it?
@@
  • Q2: How does the landscape vary across different algorithm design tasks and LLMs?
  • Q3: How do the search framework and its settings influence the landscape, and how is this connected to performance?
Our contributions are:
@@
  • We frame LLM-assisted Algorithm Search (LAS) as an optimization problem over algorithm space, where each algorithm is often represented as code (code implementation) and/or thought (algorithm idea description).
@@
  • We present a graph-based approach to visualize and analyze the fitness landscape to enhance our understanding of this significant but rarely studied topic. Each node represents one algorithm in the search process, and the edges are the connections between algorithms during the search.
@@
  • We conduct a comprehensive evaluation across 6 distinct algorithm design tasks and 6 different LLMs. We find the fitness landscape of LLM-assisted algorithm search is rugged and usually multi-modal with distinct patterns and characteristics across different algorithm design tasks and LLMs, providing insights for designing more effective LAS methods.
@@ II Formulation of LLM-assisted Algorithm Search
We formulate the LLM-assisted automatic Algorithm Search (LAS) as follows. Consider a target algorithm design task $T$ (e.g., design a heuristic for the Traveling Salesman Problem) with a set of instances $I$, where $I$ represents a practical distribution of real-world scenarios. The objective is to generate an optimal algorithm $a^{*}$ within an algorithm space $A^{S}$ that minimizes (without loss of generality) a performance metric $F(a,I)$, defined as the sum of candidate performance scores $f(a,i)$ across all instances $i\in I$. Mathematically, this is expressed as:
$F(a,I)=\sum_{i\in I}f(a,i),\quad a\in A^{S}.$
@@
$a^{*}=\underset{a\in A^{S}}{\arg\min}\,F(a,I).$
The optimization problem in algorithm space is usually tackled by an iterative search pipeline [7, 8], where each algorithm is represented as code and/or text. Starting from an initial population of algorithms $P_{0}$, the process iteratively conducts three key steps—generation, evaluation, and update—until a stopping criterion is met. A general description of the pipeline is as follows:
Step 1: Initialization The initial population of algorithms $P_{0}$ is generated either by prompting LLMs or by using existing hand-crafted algorithms. While some implementations (e.g., gradient-based search and neighborhood methods) use a population size of one, most LAS works employing evolutionary frameworks typically maintain larger population sizes.
Step 2: Iterative Process For the current algorithm population $P_{i}$, the iterative process consists of three sub-steps:
  • Step 2.1: Generation Pre-trained LLMs are prompted based on instructions and existing parent candidates $a_{p}$, which are selected from the current population $P_{i}$. This can be represented as $a_{o}=LLM(Prompt(a_{p}))$, where $a_{p}$ and $a_{o}$ denote the parent and offspring candidates, respectively, and $LLM$ represents the LLM inference process guided by instructions. $Prompt()$ is the instruction used for LLMs.
  • Step 2.2: Evaluation The fitness $F(a,I)$ of each offspring $a$ candidate is computed over a set of representative instances $I$. The average performance is usually used as fitness.
  • Step 2.3: Update The population is updated to $P_{t+1}$ by incorporating new candidates $a_{o}$ into $P_{t}$, based on some population management strategies (e.g., elitism, generational replacement).
Step 3: Termination The search process terminates when a stopping criterion is met. This criterion could be a maximum number of iterations, a maximum number of LLM requests, a performance threshold, or other problem-specific conditions. Upon termination, the best-performing algorithm $\hat{a}$, along with any other relevant information, is recorded.
Input: Initial population $P_{0}$, Instructions for LLMs, Stopping criterion
Output: Best-performing algorithm $a^{*}$ and relevant information
@@
$P\leftarrow P_{0}$ ;
  // Initialize population with $P_{0}$ (from LLMs or existing algorithms)
@@
while not Stopping criterion do
       $a_{p}\leftarrow P$ ;
        // Select parent from $P$
       $a_{o}\leftarrow LLM(Prompt(a_{p}))$ ;
        // Generate offspring using LLMs and parent $a_{p}$
       $F(a_{o},I)\leftarrow Evaluate(a_{o},I)$ ;
        // Evaluate on instances $I$
       $P\leftarrow UpdatePopulation(P,a_{o})$ ;
        // Update population with new offspring
@@
Algorithm 1 LAS Pipeline
Figure 1: Three distinct algorithm representations are commonly used in automated algorithm search methods. i) For continuous/discrete space, many existing optimization techniques can effectively handle the optimization of fitness landscapes. ii) In the case of tree-based/structured representations, methods such as genetic programming are often employed. iii) Code-based/text-based representations, on the other hand, offer greater flexibility and power. Recent advancements in LLMs have enabled direct operation and search within this complex space. However, it poses significant challenges in understanding the corresponding fitness landscape.
@@ II-A Algorithm Representation
Prior to the advent of LLMs, most existing works [17, 18, 19] on automated algorithm design can be classified into two categories based on their approaches for defining the algorithm search space $a\in A^{S}$:
@@
    Regular continuous/discrete space: $A^{S}$ is parameterized as a (continuous, discrete, or mixed) parameter space [20], which can be solved as a classic optimization problem.
@@
    Tree-based/structure-based space: Each algorithm $a$ is expressed as a tree or graph structure, and search methods such as genetic programming can be used to find a better algorithm [21, 22, 23].
@@
Unlike traditional approaches, LLM-assisted algorithm search represents each algorithm $a$ using a code implementation [24] and/or linguistic description [2]. This representation provides a richer, more expressive search space than conventional methods. However, because this space is not rigidly structured, it is challenging to understand the landscape and navigate it effectively. Figure 1 contrasts three distinct algorithm representations: (1) regular continuous/discrete space, (2) tree-based/structure-based space, and (3) code-based/text-based space.
@@ II-B Algorithm Generation
$a_{o}=\text{LLM}(\text{Prompt}(a_{p})).$
Here, the offspring algorithm $a_{o}$ is produced by prompting the LLM with specific instructions $\text{Prompt}()$, using parent algorithms $a_{p}$ as input. In a single-point iterative search, one offspring is generated from a single parent in each iteration. In contrast, population-based search maintains and evolves a population of algorithms, where both the number of parent and offspring algorithms typically exceeds one. Both the prompt strategies employed and the capabilities of the LLMs will influence the algorithm generation and thus may lead to different landscapes and performance.
@@ II-C Evaluation
III Graph-Based Fitness Landscape
@@
  • $S$ is the set of potential solutions, representing the search space;
  • $NS$ is the neighborhood structure, a function that assigns to each solution $s\in S$ a set of neighbors $N(s)$;
@@
In the context of LLM-assisted algorithm search, the search space $S$ comprises candidate algorithms generated by LLMs. The objective function $f$ is defined by specific performance metrics relevant to the task at hand. Since algorithms lack a well-defined distance metric or neighbourhood structure [12], the neighbourhood is instead characterized by correlations between algorithms. Following the approach in [26], we define the neighborhood of an algorithm $a_{i}$ relative to another algorithm $a_{j}$ as follows: $a_{i}\in NS(a_{j})$ if $a_{i}$ is generated from $a_{j}$ (i.e., $a_{i}=LLM(Prompt(a_{j}))$). This formulation enables the introduction of a graph-based representation for illustrating the fitness landscape.
@@
  • Nodes: Each node in the graph represents an algorithm $a_{i}$ generated during the search process. The set of all nodes in the graph is denoted by $N$, with its size represented as $n$.
  • Edges: Edges in the graph are directed and represent the genetic relationship between a parent algorithm and its offspring. Specifically, an edge exists from algorithm $a_{i}$ to algorithm $a_{j}$ if $a_{j}$ is directly derived from $a_{i}$ through the application of LLM-based search operators, such as $a_{j}=LLM(Prompt(a_{i}))$. Here, $Prompt()$ can involve LLM-based crossover [1], mutation [1], or other prompt strategies [2, 4]. Edges are weighted based on the frequency or likelihood of the transition, where the weight corresponds to the number of times the transition (i.e., the generation of $a_{j}$ from $a_{i}$) occurs during the search process.
  • Graph: The fitness landscape is represented as a graph $G=(N,E)$, where $N$ is the set of nodes corresponding to algorithms, and $E$ is the set of directed edges denoting parent-offspring relationships. Edge weights reflect transition probabilities between nodes. We consider an undirected graph, where there is no direction between $a_{i}$ and $a_{j}$ if they are connected.
Figure 2 illustrates a two-dimensional graph representing a fitness landscape, where nodes (colored by fitness value) denote algorithms and edges represent their connections. Each algorithm is a constructive heuristic for the traveling salesman problem, taking inputs like node IDs and a distance matrix to determine the next node in the solution.
IV Experiments
IV-A Date Collection
We demonstrate it on EoH [2, 1], a commonly used LLM-assisted algorithm search framework based on evolutionary search. We run all the experiments on LLM4AD  [^2] https://github.com/Optima-CityU/llm4ad Platform [27] across different tasks and LLMs. The experimental settings are as follows:
Evolutionary Search Settings
The population size of EoH is set to be 20 and the maximum number of candidate evaluations is 2,000 for each run.
@@ IV-A Date Collection
We investigate six algorithm design tasks including four heuristic algorithm design tasks for combinatorial optimization problems (Online Bin Packing (OBP[3], Traveling Salesman Problem (TSP[1], Capacitated Vehicle Routing Problem (CVRP[27], and Vehicle Routing Problem with Time Windows (VRPTW)) [27], one agent design task for the car mountain system (Car[27, 28] and one symbolic regression task aimed at approximating the curve of bacterial growth (Bact[29]. Our goal is to explore tasks across diverse domains, and we consider the symbolic regression task as a form of simple algorithm design.
@@
We tested a diverse set of commonly used LLMs of different sizes and capabilities, including GPT-3.5-Turbo, GPT-4o-Mini, DeepSeek-V3, DeepSeek-Chat, Qwen2-Turbo, and GLM-3-Turbo.
We collect all the algorithms and record the parent-offspring algorithm connections during evolutionary search. We will have one graph for each run for fitness landscape analysis.
IV-B Metrics
The following metrics were employed to evaluate the features and characteristics of the graph:
  • Density: This metric quantifies the density of connections in a graph, defined as the ratio of the number of edges to the maximum possible number of edges. It ranges from 0 (no edges) to 1 (fully connected graph). Density provides insights into the sparsity or compactness of the network, with higher values indicating greater connectivity.
  • Average Degree: The average degree represents the mean number of edges per node in the graph. It is computed by summing the degrees of all nodes and dividing by the total number of nodes. This metric reflects the overall connectivity of the graph, where a higher average degree signifies a more interconnected network.
  • Clustering Coefficient: This metric captures the tendency of nodes to form tightly-knit clusters. It is defined as the ratio of the number of triangles (closed triplets) to the number of possible triplets in the graph. The clustering coefficient ranges from 0 (no clustering) to 1 (fully clustered), indicating the presence of local community structures within the network.
IV-C Landscape on Different Tasks
In this section, we illustrate and analyze the fitness landscapes across different algorithm design tasks. Figure 3 illustrates the 2D and 3D visualizations of the fitness landscapes for five distinct tasks. In the 3D visualizations, the z-axis represents fitness values, where smaller (darker) values indicate better performance. The red node denotes the algorithm with the lowest fitness value. Figure 4 shows the probability distribution of algorithm fitness values, which are normalized to the range [0, 1].

### v3

@@ Abstract
Using Large Language Models (LLMs) in an evolutionary or other iterative search framework have demonstrated significant potential in automated algorithm design. However, the underlying fitness landscape, which is critical for understanding its search behavior, remains underexplored. In this paper, we illustrate and analyze the fitness landscape of LLM-assisted Algorithm Search (LAS) using a graph-based approach, where nodes represent algorithms and edges denote transitions between them. We conduct extensive evaluations across six algorithm design tasks and six commonly-used LLMs. Our findings reveal that LAS landscapes are highly multimodal and rugged, particularly in combinatorial optimization tasks, with distinct structural variations across tasks and LLMs. Moreover, we adopt four different methods for algorithm similarity measurement and study their correlations to algorithm performance and operator behaviour. These insights not only deepen our understanding of LAS landscapes but also provide practical insights for designing more effective LAS methods.
@@ I Introduction
Large Language Models (LLMs) assisted evolutionary or iterative search methods have been used for automated algorithm design with great success [1] [2] [3] [4] [5] [6] [7]. These methods cast an algorithm design task as an optimization problem and try to find an optimal algorithm in an algorithm space in an iterative manner. Leveraging LLM capabilities in language comprehension, reasoning and code generation, they are able to conduct effective and efficient search in an algorithm space [8]. However, there is lack of good understanding on why and when these methods work or not. Traditional math analysis can be hardly used to study these methods due to the complexity of the algorithm space [9] and the black-box nature of LLMs [10].
Fitness landscape analysis [11, 12, 13, 14, 15, 16, 17, 18, 19] is widely used for investigating behaviors of search methods, particularly, in the area of heuristics. Very recently, [20] has made a first attempt to study the landscape of LLM-assisted algorithm design methods. They introduced Abstract Syntax Tree (AST)-based Code Evolution Graphs (CEGs) and investigated the relationship between candidate algorithms  [^1] For the sake of clarity, we call an algorithm in the algorithm space a candidate algorithm (i.e., an element in the algorithm space) and their characters in feature space. However, CEGs focus on the structural properties of the algorithms found during the search, while fitness landscape analysis is less discussed.
@@
In this paper, we focus on the fitness landscape of LLM-assisted Algorithm Search (LAS). Leveraging a graph-based approach, we visualize and analyze the fitness landscape, providing deeper insights into the search dynamics of LAS. Specifically, we try to answer the following research questions:
@@
  • Q1: What does the fitness landscape of LLM-driven algorithm search look like, and what features (e.g., plateaus, ruggedness) can be used to characterize it?
@@
  • Q2: How does the landscape vary across different algorithm design tasks, LLMs, and settings?
  • Q3: How can we measure the similarity in algorithm space, and what are the correlations of these similarity measurements and algorithm behaviour?
Our major contributions are:
@@
  • We cast LLM-assisted Algorithm Search (LAS) as an optimization problem over the algorithm space, where each candidate algorithm is represented as code (code implementation) and/or thought (algorithm idea description) as in  [1, 5].
@@
  • We use a graph-based approach to visualize and analyze the fitness landscape to enhance our understanding of its structure. Each node represents one candidate algorithm generated in the search process, and the edges are the connections between algorithms during the search.
@@
  • We conduct a comprehensive evaluation across six distinct algorithm design tasks and six different LLMs. We find the fitness landscape of LLM-assisted algorithm search is usually multi-modal with distinct patterns and characteristics across different algorithm design tasks and LLMs.
  • We evaluate four code-based similarity metrics for measuring distances between algorithms in code representation. While we demonstrate correlations between code similarities and performance, validating that different evolutionary operators provide varying degrees of exploration in the algorithm space, our results also reveal high variability in these distance measurements, suggesting the need for new similarity metrics specifically tailored for LAS.
@@ II Formulation of LLM-assisted Algorithm Search
We formulate the LLM-assisted automatic Algorithm Search (LAS) as follows. Consider a target algorithm design task $T$ (e.g., design a heuristic for the Traveling Salesman Problem (TSP)) with a set of instances $I$, The goal is to find an optimal algorithm $a^{*}$ in an algorithm space $A^{S}$ that minimizes a performance metric $F(a,I)$, which is defined as the average of performance scores of $a$ across all instances $i\in I$. Without loss of generality, in this paper, we consider the minimization problem. Mathematically, it can be expressed as:
$F(a,I)=\frac{1}{|I|}\sum_{i\in I}f(a,i),\quad a\in A^{S}.$ (1)
@@
$a^{*}=\underset{a\in A^{S}}{\arg\min}\,F(a,I).$ (2)
The optimization problem in the algorithm space $A^{S}$ can be solved by evolutionary search [8, 9], where each algorithm is represented as code and/or text. Starting from an initial population of algorithms $P_{0}$, the evolutionary search iteratively conducts three key steps: generation, evaluation, and update, until a stopping criterion is met. It works as follows:
Step 1: Initialization The initial population of candidate algorithms $P_{0}$ is either produced by prompting LLMs or selected from existing hand-crafted algorithms.
Step 2: Iterative Process To produce $P_{t+1}$ from $P_{t}$, it does the following:
  • Step 2.1: Generation LLMs are prompted using instructions and existing parent candidates $a_{p}$, which are selected from the current population $P_{i}$. This can be represented as
    $a_{o}=LLM(Prompt(a_{p})),$
    where $a_{p}$ and $a_{o}$ denote the parent and offspring candidates, respectively. It should be pointed out that $a_{p}$ and/or $a_{o}$ can be multiple candidate algorithms. $LLM$ represents the LLM inference process with instructions. $Prompt()$ is the instruction used for LLMs.
  • Step 2.2: Evaluation The fitness $F(a,I)$ of $a$ is obtained by evaluating $a$ on the instance set $I$.
  • Step 2.3: Update Add some selected candidates from $a_{o}$ to $P_{t}$ and remove some candidates from $P_{t}$ to generate $P_{t+1}$.
Step 3: Termination The search process terminates when a stopping criterion is met. This criterion could be a maximum number of iterations, a maximum number of LLM requests, a performance threshold, or other problem-specific conditions. Upon termination, the best-performing algorithm $\hat{a}$, along with any other relevant information, will be the output.
Input: Instance set II, initial population generation method, selection strategy, stopping criterion
Output: Best algorithm $\hat{a}$ found
@@
Generate initial population $P_{0}$ either by prompting LLMs or selecting from existing algorithms;
Evaluate initial population;
$t\leftarrow 0$
@@
while stopping criterion not met do
 Generation: Select parent(s) $a_{p}\in P_{t}$ according to selection strategy;
   Generate offspring $a_{o}\leftarrow\text{LLM}(\text{Prompt}(a_{p}))$;
 Evaluation: Compute fitness $F(a_{o},I)\leftarrow\frac{1}{|I|}\sum_{i\in I}f(a_{o},i)$;
 Update: $P_{t+1}\leftarrow\text{UpdatePopulation}(P_{t},a_{o})$;
 $t\leftarrow t+1$;
@@
return $\hat{a}\leftarrow\underset{a\in P_{k}}{\arg\min}\,F(a,I)$.
Algorithm 1 LLM-assisted Algorithm Search (LAS)
Figure 1: Three distinct algorithm representations are commonly used in automated algorithm search methods. i) For continuous/discrete parameter repesentation, many existing optimization techniques can effectively handle the optimization of fitness landscapes. ii) In the case of tree-based/structured representations, methods such as genetic programming are often employed. iii) Code-based/text-based representations, on the other hand, offer greater flexibility and power. Recent advancements in LLMs have enabled direct operation and search within this complex space. However, it poses significant challenges in understanding the corresponding fitness landscape.
@@ II-A Algorithm Representation
Prior to the advent of LLMs, most existing works [21, 22, 23, 24] on automated algorithm design can be classified into two categories based on their approaches for defining the algorithm search space $a\in A^{S}$:
@@
    Continuous/discrete space: $A^{S}$ is parameterized as a (continuous, discrete, or mixed) parameter space [25, 26], which can be solved as a classic optimization problem.
@@
    Tree-based/structure-based space: Each algorithm $a$ is expressed as a tree, graph, or other structures [27], and search methods such as genetic programming can be used to find a better algorithm [28, 29, 30].
@@
Unlike traditional approaches, LAS methods usually represent each candidate algorithm $a$ using a code implementation [31] and/or linguistic description [1]. This representation is more expressive than conventional methods. However, because this space is not rigidly structured, it is challenging to understand the landscape and navigate it effectively. Fig. 1 contrasts three distinct algorithm representations: (1) regular continuous/discrete space, (2) tree-based/structure-based space, and (3) code-based/text-based space.
@@ II-B Algorithm Generation
$a_{o}=\text{LLM}(\text{Prompt}(a_{p})),$
where the offspring algorithms $a_{o}$ are produced by prompting the LLM with specific instructions $\text{Prompt}()$, using parent algorithms $a_{p}$ as input. In a single-point iterative search, one offspring is generated from a single parent in each iteration. In contrast, population-based search maintains and evolves a population of candidate algorithms, where both the number of parent and offspring algorithms typically exceeds one. Both the prompt strategies employed and the capabilities of the LLMs will influence the algorithm generation and thus may lead to different landscapes and performance.
@@ II-C Evaluation
III Graph-Based Representation of Fitness Landscape
@@
  • $S$ is the set of potential solutions;
  • NSNS is the neighborhood structure, a function that assigns to each solution $s\in S$ a set of neighbors $N(s)$;
@@
In the context of LLM-assisted algorithm search, the solutions $S$ are candidate algorithms generated by LLMs. The objective function $f$ is defined by specific performance metrics relevant to the task at hand. Since their are no well-defined distance metrics or neighbourhood structure [33, 13] in the algorithm space, the neighbourhood is instead characterized by correlations between algorithms. Similar to Search Trajectory Networks (STNs) [34] and Attractor Networks [19], where no local optima are required and nodes can be any search space location, we adapt these concepts to LLM-assisted algorithm search. We define the neighborhood of an algorithm $a_{i}$ relative to another algorithm $a_{j}$ as follows: $a_{i}\in NS(a_{j})$ if $a_{i}$ is generated from $a_{j}$ (i.e., $a_{i}=LLM(Prompt(a_{j}))$). This formulation enables us to develop a graph-based representation that effectively illustrates the fitness landscape.
@@
  • Nodes: Each node in the graph represents an algorithm $a_{i}$ generated during the search process. The set of all nodes in the graph is denoted by $N$, with its size represented as $n$. The size of each node represents the number of this algorithm generated during the search process (the same algorithm can be generated many times during the search process).
  • Edges: Edges in the graph are undirected and represent the genetic relationship between a parent algorithm and its offspring. Specifically, an edge exists from algorithm $a_{i}$ to algorithm $a_{j}$ if $a_{j}$ is directly derived from $a_{i}$ through the application of LLM-based search operators, i.e., $a_{j}=LLM(Prompt(a_{i}))$. Edges are weighted based on the frequency of the transition, where the weight corresponds to the number of times the transition (i.e., the generation of $a_{j}$ from $a_{i}$) occurs during the search process.
  • Graph: The fitness landscape is represented as a graph $G=(N,E)$, where $N$ is the set of nodes corresponding to algorithms, and $E$ is the set of edges denoting parent-offspring relationships. We consider an undirected graph, where there is no direction between $a_{i}$ and $a_{j}$ if they are connected.
Fig. 2 illustrates a two-dimensional graph representation of the fitness landscape for the traveling salesman problem. In this visualization, each node represents a distinct constructive heuristic algorithm, while edges indicate the connections between these algorithms during the LAS search process. The colour intensity of each node corresponds to its objective value, with darker shades indicating better performance (lower values). Each algorithm shown is a constructive heuristic that determines the next node selection strategy when iteratively building a TSP solution. These algorithms take several inputs: ”node c” (the current node ID), ”node d” (the destination node ID), ”nodes u” (the set of unvisited node IDs), and ”dis” (the distance matrix). Based on these inputs, each algorithm outputs the next node ID to visit, effectively defining different strategies for traversing the graph to construct a complete tour.
IV Landscape of LLM-assisted Automated Algorithm Search
IV-A Experimental Settings
We conduct Large Language Model-based Evolutionary Search (LES) on six algorithm design tasks using six widely used open-source and commercial LLMs. Our experiments employ EoH [1, 35], a standard LLM-assisted evolutionary search framework for algorithm discovery. EoH follows a simple evolutionary structure, making it suitable for analyzing fitness landscapes.
We initialize the search using EoH’s default prompt to generate initial algorithms and apply its four LLM-based evolutionary operators (E1, E2, M1, and M2) to explore new candidate algorithms. Each run maintains a population size of 20 and evaluates up to 2,000 candidate algorithms, which is the typical settings for LES [1, 4, 5].
@@ IV-A Experimental Settings
We evaluate LES across six tasks spanning multiple domains:
  • Combinatorial Optimization:
    • Online Bin Packing (OBP) [3]: Design a heuristic algorithm that assigns incoming items to bins in real time to minimize the total number of bins used.
    • Traveling Salesman Problem (TSP) [35]: Design a constructive heuristic algorithm that selects the next node given a partial route and unvisited nodes to minimize the total route length.
    • Capacitated Vehicle Routing Problem (CVRP) [36]: Design a constructive heuristic algorithm that selects the next node while satisfying capacity constraints, minimizing the total route length.
    • Vehicle Routing Problem with Time Windows (VRPTW) [36]: Design a constructive heuristic algorithm that selects the next node while satisfying both capacity and time window constraints, minimizing the total route length.
  • Agent Design:
    • Mountain Car Task (Car) [36]: A reinforcement learning task from OpenAI Gym [37] where the goal is to optimize the car’s actions to reach a target position with minimal iterations, adhering to position and velocity constraints.
  • Symbolic Regression:
    • Bacterial Growth Curve Approximation (Bact) [38]: A biology-inspired task aiming to model bacterial growth patterns by minimizing the mean squared error with respect to environmental parameters. The inclusion of symbolic regression allows us to examine algorithm/function design in a simpler, equation-based setting.
@@
We test six LLMs of varying scales and capabilities:
  • OpenAI: GPT-3.5-Turbo, GPT-4o-Mini
  • DeepSeek: DeepSeek-V3, DeepSeek-Chat
  • Other Open Models: Qwen2-Turbo, GLM-3-Turbo
All experiments are executed on the open-source LLM4AD Platform [36]. We record all generated algorithms and their parent-offspring relationships during evolution, constructing a fitness landscape graph for each run.
IV-B Landscape on Different Tasks
In this section, we illustrate and analyze the fitness landscapes across different algorithm design tasks. Fig. 3 illustrates the 2D and 3D visualizations of the fitness landscapes for five distinct tasks. In the 3D visualizations, the z-axis represents fitness values, where smaller (darker) values indicate better performance. The red node denotes the algorithm with the best value. Fig. 4 shows the probability distribution of algorithm fitness values, which are normalized to the range [0, 1].